OpenStudy (errrwut):

Link to document inside I do not understand the intermediate steps that were taken to solve for a constant of an integral.

OpenStudy (errrwut):

http://people.whitman.edu/~hundledr/courses/M367/HWSect01_4.pdf Question 1.4.2 (B) Where they solve the ODE: \[u _{xx} = -\frac{ Q }{ K_{0} } = -x \rightarrow u_{x} = -\frac{1}{2}x ^{2}+\frac{1}{6}L^{2}\] I do not understand where they got: \[+\frac{1}{6}L^{2}\] Because if you solve the integral: \[\int\limits_{0}^{L}u_{xx} dx =\int\limits_{0}^{L}(-x) dx \] Don't you just get \[-\frac{x^{2}}{2}+C_{1}\] Right?